wiki:UC20180905

Version 1 (modified by trac, 11 months ago) ( diff )

--

APO 3.5-m Users Committee Telecon, 9/05/18

Attending: Nancy Chanover (NMSU), Mark Klaene (APO), Leslie Hebb (NAPG), Misty Bentz (GSU), Alec Brown (CU), Chip Kobulnicky (UWy), Eric Bellm (UW), Mary Beth Kaiser (JHU), John Wisniewski (OU), Russet McMillan (APO), Moire Prescott (NMSU), Ben Williams (UW), Kal Kadlec (UW), Mark Whittle (UVa)


User feedback and comments from institutional representatives:

  • Wyoming - nothing to report
  • Washington - nothing to report
  • UVa - nothing to report
  • Seattle - no report
  • Oklahoma - nothing to report
  • NMSU - nothing to report
  • NAPG - nothing to report
  • JHU - nothing to report
  • Georgia State - nothing to report
  • Colorado - nothing to report

Telescope and Instruments Report:

Mark's detailed report is included below, followed by additional information discussed during today's call.

3.5-m Telescope and Instruments Highlights, 7/31/18 through 8/31/18

0) Overview

Much of the basic telescope maintenance has been accomplished. Due to metrology issues at Hextek the tertiary was never polished and was thus sent to coaters to be coated along with M2. Both mirrors had protective aluminum coating installed and are now back at APO and are in the process of being reinstalled.

1) Telescope

Final assembly, collimating, and on-sky engineering is in progress.

Full shutdown report will be available at the telecon.

2) Instruments

DIS: Upon vacuum servicing the red dewar would not reach satisfactory vacuum. New o-rings were installed again and further leak testing revealed still further issue in the electrical feed-thrus. It appears this is at the solder joint that seal the connectors themselves, not the pins. New plates and connectors are being ordered and fabricated and the exisiting plate being rehabbed at UW. This involves delicate work in the dewar and with the CCD. The blue camera will likely have similar issues in the future and we hope to wait until next year to address the blue camera.

TripleSpec is unchanged.

Agile is unchanged.

Echelle vacuum has been serviced and is now cold with excellent IOL.

NIC-FPS is operational.

ARCTIC servicing is underway as well as continued work on filter wheel repeatability.

3) ARCSAT

New guider scope has been installed and waiting for more engineering time to adjust and commission. Some data collection has been ongoing.


Additional telescope and instrument discussion, including summer shutdown overview:

The good news is that we now have M2 and M3 back from the coaters and installed on the telescope. The mirrors arrived back at APO last Friday morning and the engineering staff worked all weekend to put the telescope back together. M2 came out pretty well. As a reminder of the M3 situation, there appears to be a fundamental flaw with the standard chemical stripping solution that almost every observatory uses on their mirrors: if it is applied enough times it will etch the glass. We believe this is what happened to ARCSAT and also to M3. We therefore sent M3 to Hextek for repolishing; problems ensued and ultimately M3 was extracted from Hextek (without having been polished) and sent directly to the coaters. The coaters confirmed that there is some etching in the glass of M3. The new coating is better than what we had seen from Sunspot coatings but the effect of the etching - a fogging in some areas of the glass - is still noticeable. We are taking measurements of it now but have no direct metrics yet to predict what the impact will be. Mark saw about a 3% reduction in reflectivity in the area where cloudiness shows up vs. where it doesn’t. How we move forward next year will be discussed internally, but it is definitely good enough to be used now. We don't know whether the coating will continue to deteriorate (since previously it did so over a few month timescale as the coating continued to oxidize) - we will know more by the next UC meeting. Alec asked whether there is a change to the PSF and/or scattered light as a result of the M3 reflectivity reduction? Mark said there is definitely some scattered light, but it is much improved from where it was in June. Russet said that in the past we have not seen a change in the PSF as a result of etching.

The telescope is all back together although it has a few problems, which were not totally unexpected after coming out of a shutdown of this duration. The next tasks are to finish collimation and finish telescope characterization (pointing models etc.) before turning it over for science observations. Exactly when this happens will depend on how quickly the engineering staff can solve the minor remaining telescope issues as well as the weather (southern New Mexico is still experiencing monsoonal weather patterns).

Various instrument issues have surfaced, none of which are particularly alarming but it will take some time to work through them. DIS was serviced over the summer shutdown; some o-rings were changed out and we ended up with many leaks. Bill fixed all the o-ring leaks but still saw leaks in electrical feedthroughs. UW is making new plates with new connectors right now. We expect it to take a little more time to finish those parts; once we receive them and reassemble DIS we will need to pump on it and leak check it. There was an issue with the NICFPS filter wheel but that has been resolved. The TripleSpec guider is also fixed. The echelle is working well; it came back from servicing with reduced scattered light. There is a problem with the ARCTIC filter wheel motion but this is different because it was worked on over the summer and improvements were made. The filter positioning is now more repeatable and reliable, but sometimes it will not home filter wheel on the first try - it works every other try. This is being debugged but meanwhile it is still functional. John W requested that this issue of currently needing to rehome the filter wheel twice be added to the instrument updates web page. There are some users who are doing multicolor imaging with ARCTIC and they already have scripts written that include commands for rehoming the filter wheel. They need to know about this issue in case their scripts need to be modified. Russet said that Bill is the curator of that page but she will check with him and will add that information if she has access.


Q3 3.5-m scheduling:

The biggest impact to the remainder of the Q3 schedule is going to be the issue with DIS - it will not be ready to go this weekend even if the telescope is returned to science. We have 6 remaining OPEN/DD half nights in September but they are during bright time. We are exploring several options for paying back the displaced DIS users, including: 1) reschedule them for OPEN/DD time in September if they can use bright time, 2) look for ways to do some time payback in Q4 (we may not be able to bake it into the schedule but we would prioritize requests from displaced DIS users for Q4 OPEN/DD time), 3) examine the potential for time trades (inter- and intra-institutional trades). We ask for everyone's patience as we work through this scheduling challenge and we remind everyone that cross-institutional time trades should be brokered by the institutional scheduling reps along with Russet, Ben, and Nancy, and not by individuals contacting other individuals to request a time trade.


Q3 ARCSAT scheduling:

The new guider is installed and we are slowly making progress on commissioning it. We achieved focus but still need to get it aligned with the telescope. Meanwhile tracking is improved. There is some open time at the end of September (bright time) - anyone interested in using this time should submit a request to Russet, Ben, and Nancy.


Q4 3.5-m scheduling:

A halves in October are likely to be oversubscribed but perhaps not as much as previously thought. Nevertheless, payback time (due to DIS unavailability in September) is likely not going to be possible with October A halves. We are still waiting for one more request to come in and then the schedule will be built; we hope to distribute it next week.


Q4 ARCSAT scheduling:

The call for Q4 ARCSAT proposals will be issued after the Q3 and Q4 3.5m schedules are finalized. However, if people are interested in using any of the open time in September we welcome those requests any time.


NA2 Baffle Project

John Wilson could not make the call but provided a report to Nancy. His team has made progress over the summer with developing a SolidWorks model of the telescope and doing a backward ray trace to see how rays from the detector plane would encounter current baffle structures. The next steps are to design a new baffle that would support the larger ARCTIC FOV while simultaneously minimizing stray light and guider FOV vignetting. Then we would have that design analyzed by an optical engineering firm. Nancy will be speaking with John this month about firming up a timetable for these activities.


KOSMOS update

KOSMOS is at the UW lab, unpacked, and assembled. Now the computer is connected and talking to the instrument. The next step is to work on getting the dewar cooled and under vacuum.


TEWG Update

The Training and Education Working Group has completed their data gathering and associated discussions and is now in the process of synthesizing the information into a coherent document. The end result will be a white paper that describes current uses of the 3.5m for training and education, as well as suggestions for best practices and future approaches. The hope is that this white paper will be completed within the next couple of months.


Response to Call for White Papers for 2020 Decadal Survey

Nancy contacted several observatory directors regarding a possible white paper describing the roles that 3-4m class telescopes can play in observational astronomy in the next decade. The response she received has been mixed (either positive or no response), so Nancy will proceed with gathering input from those who are interested in participating.


ARC Board of Governors Meeting

The ARC BoG meeting is in two months, on Monday November 5. As has been done previously, we would like to showcase science highlights from the past year from all ARC partner and leasing institutions. Nancy will likely not have time to verbally explain each highlight during her presentation, but she would like to have a comprehensive packet that she can provide to the Board members for their reference. It is extremely important for the Board members to see that their investment in the 3.5m continues to benefit their institutions. To that end, Nancy is requesting that these science highlights be sent to her no later than October 15. The specifics are as follows:

  • deadline: October 15
  • format: Powerpoint, Keynote, or PDF
  • length: one slide
  • content:
    • name(s) and institution of researcher(s)
    • title of project
    • one or more graphic showing the result (image, spectrum)
    • a text block explaining the result
    • paper citation if there is one

Individuals can send their slides to Nancy directly, or an institution can designate someone (e.g. the Users Committee rep!) to collect the slides and send them to Nancy as a single packet.


ACTION ITEMS:

Action items from previous meetings:

  • Nancy will provide email updates to the Users Committee and Institutional Schedulers when there are updates on the scheduled return to science. STATUS: CLOSED. This was done today during the UC meeting and Russet has been in contact with individuals whose scheduled time may be impacted by the return to science.
  • Users Committee reps are asked to notify their users who are scheduled to observe in September of the summer shutdown schedule slip. STATUS: CLOSED.
  • Users Committee reps are asked to communicate to their faculty that if anyone is planning to propose a class training trip to the site for Q4 2018 they should notify Nancy, Ben, and Russet ASAP (before the proposal deadline!). STATUS: CLOSED.
  • Nancy will check with Russet on the status of the ARCTIC dark current issue. STATUS: CLOSED. This was briefly discussed. Russet agreed that the calibration section of the ARCTIC users manual needs to be updated but those efforts are on hold until the Q3 and Q4 schedules are complete. ARCTIC calibrations should be done as one would do for standard photon detector behavior. The darks should scale with exposure time but should not be applied to exposures less than 30 sec. For flats (for both narrowband and broadband imaging), sky flats are best, and dark sky flats are even better if one can dither the detector to create these. There is a known fringing issue that is a problem at some wavelengths (broadband).
  • Nancy will connect John Bally and Joanne Hughes Clark re: ARCTIC distortion maps. STATUS: CLOSED. Introduction email sent 8/6/18. Email discussion among several interested parties is ongoing.

New action items from this meeting:

  • Bill and/or Russet will update the Instrument Status web page to reflect the current status of the ARCTIC filter wheel positioning (need to rehome twice).
  • All member institutions and leasing partners should collect science highlight slides and send them to Nancy no later than October 15.
  • Russet will update the calibration section of the ARCTIC users manual once the Q3/Q4 scheduling tasks are completed.

Next meeting:

The next Users Committee telecon will be on Monday October 1 at 10:30 am MDT.

Note: See TracWiki for help on using the wiki.